# SWIFT:A Scalable lightWeight Infrastructure for Fine-Tuning (work in progress)

### v3

@@ Abstract.
Recent development in Large Language Models (LLMs) and Multi-modal Large Language Models (MLLMs) have leverage Attention-based Transformer architectures and achieved superior performance and generalization capabilities. They have since covered extensive areas of traditional learning tasks. For instance, text-based tasks such as text-classification and sequence-labeling, as well as multi-modal tasks like Visual Question Answering (VQA) and Optical Character Recognition (OCR), which were previously addressed using different models, can now be tackled based on one foundation model. Consequently, the training and lightweight fine-tuning of LLMs and MLLMs, especially those based on Transformer architecture, has become particularly important. In recognition of these overwhelming needs, we develop SWIFT, a customizable one-stop infrastructure for large models. With support of over LLMs and MLLMs, SWIFT stands as the open-source framework that provide the most comprehensive support for fine-tuning large models. In particular, it is the first training framework that provides systematic support for MLLMs. In addition to the core functionalities of fine-tuning, SWIFT also integrates post-training processes such as inference, evaluation, and model quantization, to facilitate fast adoptions of large models in various application scenarios. With a systematic integration of various training techniques, SWIFT offers helpful utilities such as benchmark comparisons among different training techniques for large models. For fine-tuning models specialized in agent framework, we show that notable improvements on the ToolBench leader-board can be achieved by training with customized dataset on SWIFT, with an increase of 5.2%-21.8% in the Act.EM metric over various baseline models, a reduction in hallucination by 1.6%-14.1%, and an average performance improvement of 8%-17%.
@@ 1. Introduction
In the past few years, Transformer (NIPS2017_3f5ee243, ) have been widely recognized as one of the dominant architectures for large models. In the early stages, encoder-only structures are utilized to accomplish tasks such as text-classification and sequence-labeling, with models like BERT(devlin2018bert, ) serving as typical examples. Conversely, encoder-decoder and decoder-only structures were used mostly for text generation tasks. In comparison, vision models often adopts ResNet architecture to handle tasks like Visual Question Answering (VQA), Object Detection, and Image Segmentation. These early approaches to deep learning tasks were characterized by the use of distinct model structures for different tasks.
The abundance in computational power and structured training data has brought out potentials of Transformer-based models, as industry begins to overtake models tuned for single-task. This shift positions Transformer as the preferred architectures for open-domain applications. Most notable examples include the GPT models(NEURIPS2020_1457c0d6, ; radford2019language, ) from OpenAI, as well as M6 (lin2021m6, ) and OFA (pmlr-v162-wang22al, ) models. Such progress highlighted the feasibility of using a single model to address multiple closed-domain tasks. Consequently, the paradigm of leveraging large-scale foundation models for generative tasks has emerged as the new standard for addressing multiple tasks including text-classification and sequence-labeling. The attention mechanism has also gained traction in addressing distinct multi-modal tasks with one foundation model. The release of Qwen-VL (bai2023qwen, ), GLM4-V (glm2024chatglm, ), InternVL (chen2023internvl, ), and DiT models (peebles2023scalablediffusionmodelstransformers, ) , can all satisfy to this trend. These foundation models exhibit robust capabilities in open-domain image-text and video-text question answering, as well as in image generation. They have also shown potentials for recognizing detailed image information and performing bounding-box annotation, achieving results comparable to previous closed-domain solutions.
@@
Given the large number of parameters and high memory consumption of large models, their out-of-box training has become as a significant bottleneck in the proliferation of AI. Early solutions, such as Prefix Tuning (li2021prefixtuning, ), Prompt Tuning (lester-etal-2021-power, ), and P-Tuning (liu2021gpt, ; DBLP:journals/corr/abs-2110-07602, ) open the chapter of resource-efficient training, but they can suffer from “knowledge forgetting” – a phenomenon description the scenario where fine-tuned LLMs may lost it general capacities from the foundation models. The introduction of LoRA (hu2022lora, ) shows the potentials of reduces memory consumption during training significantly comparing to full-parameter training, without loss of generality in models. This allows developers to embark on efficient-training on domain data using hardware that is much easier to access. Subsequently, more similar techniques were introduced, such as the enhancement algorithm rsLoRA (kalajdzievski2023rankstabilizationscalingfactor, ), DoRA (liu2024doraweightdecomposedlowrankadaptation, ), PISSA (meng2024pissaprincipalsingularvalues, ), OLoRA (buyukakyuz2024oloraorthonormallowrankadaptation, ), LoRA+ (hayou2024loraefficientlowrank, ), and the LLaMA-Pro (wu2024llamaproprogressivellama, ) have proliferated, providing an array of new techniques for efficient fine-tuning. In recognition of the vast differences among these different techniques, efforts begin to emerge to unify training interfaces. For example, Hugging Face has introduced the PEFT 444GitHub: https://github.com/huggingface/peft that specializes in collecting and standardizing interfaces for efficient fine-tuning algorithms.
@@
In addition to lightweight training techniques based on additional structures such as LoRA, quantization stands as another solution for reducing memory consumption during training. Typically, LLMs use 16-bit half-precision formats, such as float16 and bfloat16, for both inference and training. By reducing the tensor types to 8-bit or 4-bit, the same model can be loaded with less memory. It is even possible to run the model with 1-bit or 1.5-bit precision; this approach is collectively known as quantization. For instance, BitsAndBytes (dettmers2022llmint8, ) employs segmented quantization with outlier thresholds, AutoGPTQ (frantar2023gptqaccurateposttrainingquantization, ) performs Taylor series decomposition on parameter matrices and uses the Hessian matrix to evaluate parameter importance, and AWQ (lin2024awqactivationawareweightquantization, ) evaluates parameter significance and applies scaling factors for quantization. Due to the complexity of quantization techniques and their poor adaptability to different models, the Hugging Face community has introduced the Optimum library555GitHub: https://github.com/huggingface/optimum as a unified implementation for various quantization methods. Nevertheless, the task of LLM training and fine-tuning remains formidable for most developers, as the aforementioned solutions only cover support of a limited number of models and techniques. In particular, support for newer models and techniques are often lacking in existing solutions. Furthermore, to ensure that the trained models can be effectively deployed, the post-training processes, such as inference and evaluation, are also steps in utilization of trained models. To address this, we introduced SWIFT 666GitHub:https://github.com/modelscope/swift, an open-source framework targeted at facilitating lightweight training of large models, which also incorporates comprehensive functionality for post-training processes. SWIFT assists developers to perform training and inference operations with minimal learning overhead. By streamlining various technical components, sourced or self-developed, in a unified way, SWIFT is tasked to enable efficient training and development pipelines of large models.
@@
We introduce SWIFT, a training framework compatible with the general model standards of the Transformers library. SWIFT integrates libraries such as PEFT and Optimum, enabling pre-training, fine-tuning, and human alignment for LLM and MLLM models. In addition quantization training (QLoRA (dettmers2023qloraefficientfinetuningquantized, )) is supported as well. Today SWIFT supports over 300 LLM models and over 50 MLLM models, encompassing all major open-source models. It also comes with support of over 150 pure text and multi-modal datasets.
Other than the standard Attention structures, training and inference for model structures such as Mamba (gu2024mambalineartimesequencemodeling, ) model are also supported in Swift. Furthermore, training with Megatron (shoeybi2020megatronlmtrainingmultibillionparameter, ) structured models is supported as well which facilitates large-scale parallel pre-training across multiple nodes and GPUs to be performed with SWIFT.
@@
| MLLM-PRETRAIN | 3 models | 50+ models | ||||
| MLLM-SFT | 3 models | 50+ models | ||||
| MLLM-RLHF | 3 models | 50+ models | ||||
@@ 2. Related Works
LLaMA-Factory(zheng2024llamafactoryunifiedefficientfinetuning, ) is a versatile, all-in-one large model training framework. This framework is fully compatible with the Hugging Face model ecosystem. Additionally, it supports a WEB-UI based on Gradio, further reducing the cost of usage. LLaMA-Factory supports the pre-training, fine-tuning, and human alignment of over 100 text LLMs. It also facilitates the training of some of them multi-modal models such as LLaVA, PaliGemma, and YI-VL. In terms of evaluation capabilities, it supports the evaluation processes for the CEVAL, MMLU, and CMMLU datasets and enables inference and deployment workflows based on vLLM.
Firefly777GitHub:https://github.com/yangjianxin1/Firefly leverages the transformers training ecosystem (Trainer/PEFT, etc.). Remarkably, it explores training datasets and creates several popular datasets for NLP training, such as firefly-train-1.1M, moss-003-sft-data, and ultrachat. They have also utilized these dataset for lightweight training on various models, including firefly-mixtral-8x7b, which has outperformed Yi-34B-Chat on multiple leaderboards.
FastChat(zheng2023judging, ) is a comprehensive training and inference framework. This framework is equipped with capabilities for model training, inference, deployment, and evaluation. FastChat has leveraged Transformers and PEFT for training, and supports models such as LLaMA, Vicuna, and T5. It supports lightweight fine-tuning using techniques such as LoRA, QLoRA, and XFormers888GitHub: https://github.com/facebookresearch/xformers. For deployment, FastChat supports inference acceleration frameworks like vLLM, SGLang999https://github.com/sgl-project/sglang, and LightLLM101010https://github.com/ModelTC/lightllm. FastChat places a focus on model inference and deployment, with relatively limited training support.
Axolotl111111https://github.com/axolotl-ai-cloud/axolotl employs training component libraries such as TRL, PEFT, and Transformers. This framework has extended training capabilities, including the encapsulation of the mambassm121212GitHub: https://github.com/state-spaces/mamba library, thereby enhancing the ability to train these models using the transformers ecosystem. Axolotl supports the LoRA and QLoRA training of various models across multiple series, including LLaMA, Mistral, Qwen, and Phi, and also supports inference and merge-LoRA operations.
@@ 3.1. Design of Training Architecture
Specifically, in the direction of pre-training, SWIFT supports Megatron architecture models. Particularly in the CPT scenario, SWIFT first converts the checkpoints of the transformers architecture into the checkpoints of the Megatron architecture and then continues pre-training the model using various parallel methods of Megatron. After training, the checkpoints can be converted back to the format supported by transformers. SWIFT utilizes the PAI-Megatron-Patch framework 131313GitHub: https://github.com/alibaba/Pai-Megatron-Patch for supporting Megatron. The conversion process of checkpoints is supported through the export module in SWIFT.
In the training component, new optimizers such as GaLore and Q-GaLore are integrated, making them readily available for use during training. To further alleviate training pressure, SWIFT supports sequence parallelism technology (2023xtuner, ), which distributes sequences across different processes under DDP conditions, thereby reducing the memory consumption for long-sequence training.
@@
SWIFT provides three commands for different tasks: ‘pt‘ for pre-training, ‘sft‘ for fine-tuning, and ‘rlhf‘ for RLHF. These commands are consistent for both pure text models and multi-modal models. For dataset selection, SWIFT supports the use of the ‘–dataset‘ option to directly use pure text and multi-modal datasets, and it also supports referencing local training files.
@@
For lightweight training, SWIFT supports QLoRA training methods. The quantization methods available include BNB, HQQ (badri2023hqq, ), EETQ (gordon2023eptqenhancedposttrainingquantization, ), AWQ, GPTQ, and AQLM (egiazarian2024extremecompressionlargelanguage, ). In terms of model support, we facilitate the training processes for over 300 NLP models and more than 50 multi-modal models. Specifically, to effectively fine-tune agents, we collaborated to create the MS-Agent dataset 141414WebPage: https://www.modelscope.cn/datasets/iic/ms_agent. This dataset is a relatively rare, high-quality Chinese fine-tuning dataset for agents. Subsequently, we updated the MSAgent-Pro dataset151515WebPage: https://www.modelscope.cn/datasets/iic/MSAgent-Pro, which adopts the ToolBench format. This dataset is very important for supervised fine-tuning (SFT) to enhance agent capabilities, and it includes the Chain of Thought (CoT) process, which significantly improves the effectiveness of multi-turn agent calls. To facilitate agent fine-tuning, SWIFT supports the ‘tools‘ field in dataset formats and allows fine-tuning training using different prompt formats (e.g., ToolBench (qin2023toolllm, ) format, ReACT (yao2023reactsynergizingreasoningacting, ) format, or other formats defined by model templates). This can be seen in the SWIFT standard dataset definitions.

### v4

@@ Abstract.
Recent development in Large Language Models (LLMs) and Multi-modal Large Language Models (MLLMs) have leveraged Attention-based Transformer architectures and achieved superior performance and generalization capabilities. They have since covered extensive areas of traditional learning tasks. For instance, text-based tasks such as text classification and sequence labeling, as well as multi-modal tasks like Visual Question Answering (VQA) and Optical Character Recognition (OCR), which were previously addressed using different models, can now be tackled based on one foundation model. Consequently, the training and lightweight fine-tuning of LLMs and MLLMs, especially those based on the Transformer architecture, have become particularly important. In recognition of these overwhelming needs, we develop SWIFT, a customizable one-stop infrastructure for large models. With support of more than LLMs, MLLMs, and almost 200 Megatron models, SWIFT stands as the open-source framework that provides the most comprehensive support for fine-tuning large models. In particular, it is the first training framework that provides systematic support for MLLMs. In addition to the core functionalities of fine-tuning, SWIFT also integrates post-training processes such as inference, evaluation, and model quantization, to facilitate fast adoption of large models in various application scenarios. With a systematic integration of various training techniques, SWIFT offers helpful utilities, such as benchmark comparisons among different training techniques for large models. For fine-tuning models specialized in agent framework, we show that notable improvements on the ToolBench leaderboard can be achieved by training with customized datasets on SWIFT, with an increase of 5.2%-21.8% in the Act.EM metric over various baseline models, a reduction in hallucination by 1.6%-14.1%, and an average performance improvement of 8%-17%.
@@ 1. Introduction
In the past few years, the Transformer (NIPS2017_3f5ee243, ) has been widely recognized as one of the dominant architectures for large models. In the early stages, encoder-only structures are utilized to accomplish tasks such as text classification and sequence labeling, with models like BERT(devlin2018bert, ) serving as typical examples. Conversely, encoder-decoder and decoder-only structures were used mostly for text generation tasks. In comparison, vision models often adopts ResNet architecture to handle tasks like Visual Question Answering (VQA), Object Detection, and Image Segmentation. These early approaches to deep learning tasks were characterized by the use of distinct model structures for different tasks.
The abundance in computational power and structured training data has brought out potentials of Transformer-based models, as industry begins to overtake models tuned for single-task. This shift positions Transformer as the preferred architectures for open-domain applications. Most notable examples include the GPT models(NEURIPS2020_1457c0d6, ; radford2019language, ) from OpenAI, as well as M6 (lin2021m6, ) and OFA (pmlr-v162-wang22al, ) models. Such progress highlighted the feasibility of using a single model to address multiple closed-domain tasks. Consequently, the paradigm of leveraging large-scale foundation models for generative tasks has emerged as the new standard for addressing multiple tasks including text classification and sequence labeling. The attention mechanism has also gained traction in addressing distinct multi-modal tasks with one foundation model. The release of Qwen2.5-VL (bai2025qwen25vltechnicalreport, ), GLM4-V (glm2024chatglm, ), InternVL3 (zhu2025internvl3exploringadvancedtraining, ), and DiT models (peebles2023scalablediffusionmodelstransformers, ) , can all satisfy to this trend. These foundation models exhibit robust capabilities in open-domain image-text and video-text question answering, as well as in image generation. They have also shown potentials for recognizing detailed image information and performing bounding-box annotation, achieving results comparable to previous closed-domain solutions.
@@
Reinforced fine-tuning(RFT) has consistently served as a crucial mechanism for enhancing model performance following Supervised Fine-Tuning. The primary methodology involves utilizing data generated either by the model itself or by alternative models, which is subsequently filtered through reward models or reward functions, and then employed for continued model training. Among the commonly utilized reinforced fine-tuning approaches are Rejection sampling Fine-Tuning(yuan2023scalingrelationshiplearningmathematical, ) and distillation. Following DeepSeek’s introduction of the DeepSeek-R1(deepseekai2025deepseekr1incentivizingreasoningcapability, ) model, RLVR methods, especially represented by Generalized Reinforcement Policy Optimization (GRPO)(shao2024deepseekmathpushinglimitsmathematical, ) emerged as a significant reinforced fine-tuning technique. The principal advantage of this approach lies in its reinforcement learning characteristics, which enable substantial improvements in metrics such as Pass@K without necessitating massive SFT data (requiring only minimal data for code initialization). In response to these technological advancements, the open-source community has developed numerous RL frameworks, including Open-R1444GitHub: https://github.com/huggingface/open-r1, veRL555GitHub: https://github.com/volcengine/verl, and OpenRLHF666GitHub: https://github.com/OpenRLHF/OpenRLHF.
@@
Given the large number of parameters and high memory consumption of large models, their out-of-box training has become as a significant bottleneck in the proliferation of AI. Early solutions, such as Prefix Tuning (li2021prefixtuning, ), Prompt Tuning (lester-etal-2021-power, ), and P-Tuning (liu2021gpt, ; DBLP:journals/corr/abs-2110-07602, ) open the chapter of resource-efficient training, but they can suffer from “knowledge forgetting” – a phenomenon description the scenario where fine-tuned LLMs may lost it general capacities from the foundation models. The introduction of LoRA (hu2022lora, ) shows the potentials of reduces memory consumption during training significantly comparing to full-parameter training, without loss of generality in models. This allows developers to embark on efficient-training on domain data using hardware that is much easier to access. Subsequently, more similar techniques were introduced, such as the enhancement algorithm rsLoRA (kalajdzievski2023rankstabilizationscalingfactor, ), DoRA (liu2024doraweightdecomposedlowrankadaptation, ), PISSA (meng2024pissaprincipalsingularvalues, ), OLoRA (buyukakyuz2024oloraorthonormallowrankadaptation, ), LoRA+ (hayou2024loraefficientlowrank, ), and the LLaMA-Pro (wu2024llamaproprogressivellama, ) have proliferated, providing an array of new techniques for efficient fine-tuning. In recognition of the vast differences among these different techniques, efforts begin to emerge to unify training interfaces. For example, Hugging Face has introduced the PEFT 777GitHub: https://github.com/huggingface/peft that specializes in collecting and standardizing interfaces for efficient fine-tuning algorithms.
@@
In addition to lightweight training techniques based on additional structures such as LoRA, quantization stands as another solution for reducing memory consumption during training. Typically, LLMs use 16-bit half-precision formats, such as float16 and bfloat16, for both inference and training. By reducing the tensor types to 8-bit or 4-bit, the same model can be loaded with less memory. It is even possible to run the model with 1-bit or 1.5-bit precision; this approach is collectively known as quantization. For instance, BitsAndBytes (dettmers2022llmint8, ) employs segmented quantization with outlier thresholds, AutoGPTQ (frantar2023gptqaccurateposttrainingquantization, ) performs Taylor series decomposition on parameter matrices and uses the Hessian matrix to evaluate parameter importance, and AWQ (lin2024awqactivationawareweightquantization, ) evaluates parameter significance and applies scaling factors for quantization. Due to the complexity of quantization techniques and their poor adaptability to different models, the Hugging Face community has introduced the Optimum library888GitHub: https://github.com/huggingface/optimum as a unified implementation for various quantization methods. Nevertheless, the task of LLM training and fine-tuning remains formidable for most developers, as the aforementioned solutions only cover support of a limited number of models and techniques. In particular, support for newer models and techniques are often lacking in existing solutions. Furthermore, to ensure that the trained models can be effectively deployed, the post-training processes, such as inference and evaluation, are also steps in utilization of trained models. To address this, we introduced SWIFT 999GitHub:https://github.com/modelscope/swift, an open-source framework targeted at facilitating lightweight training of large models, which also incorporates comprehensive functionality for post-training processes. SWIFT assists developers to perform training and inference operations with minimal learning overhead. By streamlining various technical components, sourced or self-developed, in a unified way, SWIFT is tasked to enable efficient training and development pipelines of large models.
@@
We introduce SWIFT, a training framework compatible with the general model standards of the Transformers library. SWIFT integrates libraries such as PEFT and Optimum, enabling pre-training, fine-tuning, human alignment and reinforced fine-tuning for LLM and MLLM models. In addition quantization training (QLoRA (dettmers2023qloraefficientfinetuningquantized, )) is supported as well. Today SWIFT supports over 550 LLM models and over 200 MLLM models, encompassing all major open-source models. It also comes with support of over 150 pure text and multi-modal datasets.
Other than the standard Attention structures, training and inference for model structures such as Mamba (gu2024mambalineartimesequencemodeling, ) model are also supported in SWIFT. Training with Megatron (shoeybi2020megatronlmtrainingmultibillionparameter, ) structured models is supported as well which facilitates large-scale parallel pre-training across multiple nodes and GPUs to be performed with SWIFT.
@@
| MLLM-PRETRAIN | 60+ models | 20+ models | 200+ models | |||
| MLLM-SFT | 60+ models | 20+ models | 200+ models | |||
| MLLM-RLHF | 60+ models | 20+ models | 200+ models | |||
@@ 2. Related Works
LLaMA-Factory(zheng2024llamafactoryunifiedefficientfinetuning, ) is a versatile, all-in-one large model training framework. This framework is fully compatible with the Hugging Face model ecosystem. Additionally, it supports a WEB-UI based on Gradio, further reducing the cost of usage. LLaMA-Factory supports the pre-training, fine-tuning, and human alignment of over 100 text LLMs. It also facilitates the training of some of them multi-modal models such as LLaVA, PaliGemma, and Qwen-VL. In terms of evaluation capabilities, it supports the evaluation processes for the CEVAL, MMLU, and CMMLU datasets and enables inference and deployment workflows based on vLLM.
Firefly101010GitHub:https://github.com/yangjianxin1/Firefly leverages the transformers training ecosystem (Trainer/PEFT, etc.). Remarkably, it explores training datasets and creates several popular datasets for NLP training, such as firefly-train-1.1M, moss-003-sft-data, and ultrachat. They have also utilized these dataset for lightweight training on various models, including firefly-mixtral-8x7b, which has outperformed Yi-34B-Chat on multiple leaderboards.
FastChat(zheng2023judging, ) is a comprehensive training and inference framework. This framework is equipped with capabilities for model training, inference, deployment, and evaluation. FastChat has leveraged Transformers and PEFT for training, and supports models such as LLaMA, Vicuna, and T5. It supports lightweight fine-tuning using techniques such as LoRA, QLoRA, and XFormers111111GitHub: https://github.com/facebookresearch/xformers. For deployment, FastChat supports inference acceleration frameworks like vLLM, SGLang121212https://github.com/sgl-project/sglang, and LightLLM131313https://github.com/ModelTC/lightllm. FastChat places a focus on model inference and deployment, with relatively limited training support.
Axolotl141414https://github.com/axolotl-ai-cloud/axolotl employs training component libraries such as TRL, PEFT, and Transformers. This framework has extended training capabilities, including the encapsulation of the mambassm151515GitHub: https://github.com/state-spaces/mamba library, thereby enhancing the ability to train these models using the transformers ecosystem. Axolotl supports the LoRA and QLoRA training of various models across multiple series, including LLaMA, Mistral, Qwen, and Phi, and also supports inference and merge-LoRA operations.
@@ 3.1. Design of Training Architecture
Specifically, in the direction of pre-training, SWIFT supports Megatron architecture models. Particularly in the CPT scenario, SWIFT first converts the checkpoints of the transformers architecture into the checkpoints of the Megatron architecture and then continues pre-training the model using various parallel methods of Megatron. After training, the checkpoints can be converted back to the format supported by transformers.
In the training component, new optimizers such as GaLore, Q-GaLore and Muou are integrated, making them readily available for use during training. To further alleviate training pressure, SWIFT supports sequence parallelism technology (jacobs2023deepspeedulyssesoptimizationsenabling, ), which distributes sequences across different processes under DDP conditions, thereby reducing the memory consumption for long-sequence training.
Reinforced fine-tuning has gained prominence through notable works such as Rejection sampling Fine-Tuning, ReST(gulcehre2023reinforcedselftrainingrestlanguage, ) and B-STaR(zeng2025bstarmonitoringbalancingexploration, ). This type of training methodology employs the model itself for rollout generation, followed by filtering correct responses via Outcome Reward Modeling (ORM), and subsequently retraining the model with the filtered data. This process can be iterated multiple times until model performance reaches a plateau. Empirical evidence demonstrates that Reinforced fine-tuning can improve model metrics by more than 5 points average following Supervised Fine-Tuning (SFT), without inducing significant knowledge forgetting across other evaluation benchmarks.
Similarly, the process of training a smaller model using rollout data from larger Language Models (LLMs) is generally referred to as distillation. This approach also incorporates ORM for data quality assessment. Following the emergence of Process Reward Modeling (PRM), both Rejection sampling and distillation methodologies have been enhanced by incorporating process rewards for superior data filtration. Regarding sampling techniques, various approaches including standard sampling, Diverse verifier tree search (DVTS), or Monte Carlo Tree Search (MCTS) can be employed. To facilitate this process, SWIFT provides the ‘swift sample‘ command, which supports methods such as vanilla sampling, MCTS, and DVTS. This command enables data sampling using open-source models and also supports sampling via closed-source APIs (distillation). Following the sample command, training can be conducted using either the ‘swift sft‘ or ‘swift rlhf‘ commands.
Generative Reinforcement Policy Optimization (GRPO) has emerged as a crucial method for enhancing model capabilities due to its minimal data requirements and high robustness. This has led to the establishment of a training pipeline consisting of pre-training, cold-start SFT, and RLVR(GRPO). GRPO is predominantly utilized for eliciting Chain-of-Thought (CoT) reasoning capabilities. For instance, format rewards can be implemented to compel models to engage in explicit reasoning (e.g., requiring the response to include paired ‘¡think¿¡/think¿‘ tags), followed by the provision of the correct answer (within ‘¡result¿¡/result¿‘ tags). GRPO has demonstrated substantial impact in improving models’ Pass@K performance. Currently, GRPO is widely applied to enhance LLMs’ CoT capabilities, as well as multi-modal LLMs’ Visual Question Answering (VQA) logical reasoning and complex agent pathway capabilities. In response to this trend, SWIFT supports GRPO training methodologies(along with DAPO(yu2025dapoopensourcellmreinforcement, )) for both LLMs and MLLMs, and accommodates multi-round rollouts (for complex agent training pathways). SWIFT’s GRPO implementation can readily scale to training systems with over a hundred GPUs and support models of increasing parameter sizes.
@@
SWIFT provides several commands for different tasks: ‘pt‘ for pre-training, ‘sft‘ for fine-tuning, ‘rlhf‘ for RLHF, ‘sample‘ for sampling datasets. These commands are consistent for both pure text models and multi-modal models. For dataset selection, SWIFT supports the use of the ‘–dataset‘ option to directly use pure text and multi-modal datasets, and it also supports referencing local training files.
@@
For lightweight training, SWIFT supports QLoRA training methods. The quantization methods available include BNB, HQQ (badri2023hqq, ), EETQ (gordon2023eptqenhancedposttrainingquantization, ), AWQ, GPTQ, and AQLM (egiazarian2024extremecompressionlargelanguage, ). In terms of model support, we facilitate the training processes for over 300 NLP models and more than 50 multi-modal models. Specifically, to effectively fine-tune agents, we collaborated to create the MS-Agent dataset 161616WebPage: https://www.modelscope.cn/datasets/iic/ms_agent. This dataset is a relatively rare, high-quality Chinese fine-tuning dataset for agents. Subsequently, we updated the MSAgent-Pro dataset171717WebPage: https://www.modelscope.cn/datasets/iic/MSAgent-Pro, which adopts the ToolBench format. This dataset is very important for supervised fine-tuning (SFT) to enhance agent capabilities, and it includes the Chain of Thought (CoT) process, which significantly improves the effectiveness of multi-turn agent calls. To facilitate agent fine-tuning, SWIFT supports the ‘tools‘ field in dataset formats and allows fine-tuning training using different prompt formats (e.g., ToolBench (qin2023toolllm, ) format, ReACT (yao2023reactsynergizingreasoningacting, ) format, or other formats defined by model templates). This can be seen in the SWIFT standard dataset definitions.